OpenStudy (anonymous):

the length of a rectangle is 3 more than 5times the width find the perimeter is area=92m^2

OpenStudy (anonymous):

typo its meant to be if instead of is

OpenStudy (imstuck):

Area formula is L * W. We have here that the Length is 3 more than 5 times the width, so the L = 5W + 3. Multiply that by the width (W) and get this:\[A=(5W+3)(W)\]That multiplies out to \[A=5W ^{2}+3W\]We also know that A is 92, so that goes in there too:

OpenStudy (imstuck):

\[92=5W ^{2}+3W\]We have to solve that for W, and in order to do that we have to set up a quadratic equation. Like this:

OpenStudy (imstuck):

\[5w ^{2}+3w-92=0\]That needs to be factored to find w. Let's do that, then!

OpenStudy (anonymous):

(5w+23)(w-4)

OpenStudy (imstuck):

Ok,...but when you solve those for w what do you get?

OpenStudy (imstuck):

Very good, btw!

OpenStudy (anonymous):

-23/5, -4

OpenStudy (imstuck):

5w=-23 w= -23/5. OR w - 4 = 0 w = 4 We know that length cannot be a negative number so the width is 4. Now find the length. Do that by using your L = 5w+3.

OpenStudy (imstuck):

L=5(4)+3 L = 23. Perimeter formula is P = 2L + 2W. L is 23 and W is 4 so... P = 2(23) + 2(4) P = 46 + 8 P = 54 meters

OpenStudy (anonymous):

Thanks again